answer 7,8, and 9 :) pls
007 10.0 points A potters wheel moves from rest to an angu- lar speed of 0.10 rev/s in 26.9 s. ng constant angular acceleration, Assumi what is its angular acceleration in rad/s2? Answer in units of rad/s 008 10.0 points The tape in a videotape cassette has a total length 186 m and can play for 2 h. As the tape starts to play, the full reel has an outer radius of 40 mm and an inner radius of 13 mm. At some point during the play, both reels will have the same angular speed. What is this common angular speed? Answer in units of rad/s 009 10.0 points The driver of a car traveling at 14.3 m/s ap- plies the brakes and undergoes a constant deceleration of 1.67 m/s2 How many revolutions does each tire make before the car comes to a stop, assuming that the car does not skid and that the tites have radii of 0.12 m? Answer in units of rev.
Answer #1

Ans:-

Question 7

Given data:w1 = 0 rev/s ,w2 = 0.10rev/s ∆t= 26.9s

α =∆w/∆t

= w2 –w1/∆t

= (0.10-0)rev/s/26.9s

α =3.72*10^-3rev/s^2

but we have to acceleration in rad/s^2

so we know that 1rev = 2π

α =3.72*10^-3*2πrad/rev

α = 0.023rad/s^2

Question 8

L= 186m, t= 2h= 2*60= 120m=120*60= 7200s, R01 = 40mm, Ri = 13mm,

Velocity V= L/t = 186m/7200s = 0.0258m/s

Angular speed w = v/r

V= w*r

Assume each reel has the same inner radius (Ri). Find the disk area occupied by the full reel and then distribute it such that both reels have the same outer radius:

Ro1 is the initial outer radius. Ro2 is the reel balanced outer radius

A_tape_full = Pi*(Ro1^2 - Ri^2)
2*A_tape_balance = A_tape_full, because the tape is balanced among the two reels

A_tape_balance = Pi*(Ro2^2 - Ri^2)

2*Pi*(Ro2^2 - Ri^2) = Pi*(Ro1^2 - Ri^2)

Solve for Ro2:
Ro2 = 1/2*sqrt(2*Ri^2 + 2*Ro1^2)

R02 = ½ *sqrt (2*( 13*10^-3)^2 + 2*(40*10^-3)^2)

R02= 0.030

W = V/R02 = 0.0258/0.030

W = 0.86rad/s

Question 9

Given data : v= 14.3m/s, a= -1.67m/s^2 , r= 0.12m

Here we use the kinematics equation

use v^2 = u^2 + 2as to find the distance travelled

0 = 14.3^2 + 2(-1.67)s

s = 14.3^2/3.34

= 61.23m

Circumference of wheels = 2πr

= 2π (0.12) m

= 0.7536

Number of revolutions = 61.23/0.7536

= 81.25

approximately 81 revolutions